In the economy of Robberia, the monetary base is $2,650. People hold 40% of their money in the form of currency (and this 60% as bank deposits). Banks hold 15% of their deposits in reserve. In the face of this panic, the central bank wants to conduct an open-market operation to keep the money supply at its original level. Calculate, in dollars, how much the central bank needs to transact.

1) The central bank sells x worth of bonds. Please solve for x.

Let me know if you need additional information, thank you.
Answer #1

Monetary base = 2650; Currency = 0.4; Bank Deposits = 0.6; Reserve deposit = 0.15

Currency-deposit ratio = 0.4 / 0.6 = 0.67

Money Multiplier = 1 / Reserve ratio = (1+0.67) / (0.15 + 0.67) = 2.04082

Money Supply = MB * MM = 2.04082 * 2650 = 5408

When central bank wants to keep it to the original level i.e. 5408 it would want to increase the money supply by $2758 ( = 5408 - 2650)

Thus will want to buy 1351 government bonds [= 2758 / 2.04082 = 1351]